OpenStudy (sissyedgar):

Beth measured the distance from home plate to the pitcher’s mound on a baseball field. She measured the distance to be 69 ft. The actual distance from home plate to the pitcher’s mound is 60.5 ft. What is the percent error in Beth’s measurement? Round to the nearest tenth of a percent, if necessary.

OpenStudy (anonymous):

To find percent error: error=(|measurement−actual|actual)(100)

OpenStudy (anonymous):

@sissyedgar

OpenStudy (sissyedgar):

yea.. i dont get tht

OpenStudy (anonymous):

error=measured value-actual value error=69-60.5=8.5ft % error=(error/actual value)100=(8.5/60.5)100=14% ans.

OpenStudy (sissyedgar):

oooo,ok

OpenStudy (hazelluv99):

@sissyedgar go on history chat when ur done

OpenStudy (anonymous):

are there multiple choices

OpenStudy (sissyedgar):

did u track meh down?

OpenStudy (sissyedgar):

:P

OpenStudy (anonymous):

sometimes they don't answer @HazelLuv99